OpenStudy (anonymous):

On a coordinate plane, find the midpoint (M) of a segment with the following endpoints: A(-4, 13) and B(9, 0) 1) M(5, 13) 2) M(2.5, 6.5) 3) M(13, 5) 4) M(6.5, 2.5)

OpenStudy (saifoo.khan):

Do u know the midpoint formula?

OpenStudy (anonymous):

M(X,Y) X=(-4+9)/2=2.5 Y=(13+0)/2=6.5

OpenStudy (anonymous):

it is easy, you should watch your book carefully